# Modular DNA origami-based electrochemical detection of DNA and proteins

### v1

@@ Abstract
This manuscript was compiled on December 11, 2023
@@
Figure 1: Lily pad sensors can be used for the electrochemical detection of biological analytes, here a DNA single strand. (A) A flat, disk-shaped DNA origami with a square hole (49) carrying a long dsDNA linker is assembled from a mixture of long ssDNA scaffolds, ssDNA staples, 70 staples with ${}^{\prime}$-extensions for MB reporters and for analyte binding, 20 nt MB-modified DNA oligos, and long dsDNA linkers with ssDNA overhangs at both ends. Inset shows AFM of a lily pad. (B) Two types of thiol-modified ssDNA are immobilized on a template-stripped gold surface: one for tethering the lily pad origami via hybridization with one of the linker overhang sequences, and the other for analyte binding to the gold surface. Closing events occur when a DNA analyte binds both to the binding site on the origami and the binding site on the surface, enabling electron transfer between MB reporters and the gold electrode. Inset shows an exemplary square wave voltammogram of a sensor in the open state (left), and a voltammogram in the closed state after DNA analyte addition (right). The peak current is calculated as the heights of the peaks relative to the underlying baseline.
@@ Results
Figure 1 shows the design of our lily pad device. The device consists of a 100 nm diameter disk-shaped DNA origami (see methods) that is attached to a long dsDNA linker (Fig. 1A) which, depending on experiment, ranges from 300 to 3000 base pairs (bp) in length. The distal end of the linker is attached to a gold electrode through DNA hybridization: the distal end of the linker displays a single-stranded DNA (ssDNA) overhang, and binds to a complementary ssDNA that is immobilized on the gold electrode. E-DNA/E-AB sensors often use roughened gold electrodes to maximize the surface area for their small ssDNA sensing molecules(56, 57), despite the resultant structural heterogeneity of such electrodes. Here, in contrast, we use ultraflat, template-stripped gold chips(58) to minimize the distance between the entire bottom face of the origami disk and the electrode surface upon analyte binding. Our lily pad sensors use 70 MB molecules as reporters to transfer electrons to the electrode; they are attached to one side of each origami via 20 bp-long ssDNAs hybridized to overhangs on 70 modified staple strands (Fig. 1B)—this forms a curtain of MB-modified DNA strands hanging from the origami. Detection of a ssDNA analyte sequence xy is achieved through binding to two DNA sequences, ${}^{\prime}$ and ${}^{\prime}$, which are complementary to subsequences x and y of xy, respectively. Sequence ${}^{\prime}$ is positioned on the origami as an extension to a central staple, and a thiol-modified version of ${}^{\prime}$ is immobilized on the gold electrode surface via an Au-S bond. When both x and y in a DNA analyte bind to ${}^{\prime}$ and ${}^{\prime}$ on the lily pad and the underlying electrode, the conformation of the lily pad changes from the “open” to the “closed” state. Each closing event brings the 70 MB curtain into proximity with the gold surface, facilitating electron transfer between MB and electrode and increasing voltammetric peak currents (Fig. 1B, inset).
@@
Figure 2: Lily pad closing events can be monitored via voltammetry in real-time. (A) 500 pM of each of two different ssDNA analytes, with 28-bp sequences xy (blue) and yx (orange), were added at time $t=0$ at ${}^{\circ}$. The difference in kinetics for xy and yx binding can be understood in terms of difference spacing between the origami and gold surface in their respective bound states (bottom left). For yx to bind (orange box), the origami must get much closer to the surface than in the case of xy (blue box), incurring an increased entropy penalty. Binding of yx may also be sterically hindered by the 20 bp long MB reporters that comprise a curtain projecting down from the origami (top left). (B) The length of the linker tethering the lily pad to the surface has little effect on sensor response: linker lengths of 332, 1037, and 2983 bp all show comparable behavior. (C) The functional form of sensor response varies with the concentration of ssDNA analyte. At analyte concentrations up to 500 picomolar (yellow diamonds), sensor response increases monotonically. At nanomolar concentrations (green circles), the system shows saturation effects, and the current decreases after one hour, potentially due to excess free analyte in solution reopening previously-closed lily pads via DNA strand displacement. ssDNA analyte xy added at $t=0$.
@@
Because of the lily pad’s modularity, it is trivial to modify into a detector for multimeric proteins with multiple identical binding sites. As a proof-of-principle (Fig. 3), we prepared lily pad sensors for detection of a model multimeric protein, streptavidin. We achieved this by simply adding two biotinylated adaptor strands to our basic DNA-sensing chips (Fig. 3A): the first adaptor, sequence x, was complementary to the 14 nt ssDNA tail (${}^{\prime}$) on the origami; the second, sequence y, was complementary to the 14 nt oligos (${}^{\prime}$) on the surface. Thus chips with the lily pads used for ssDNA detection were incubated with the biotin-modified adaptors to obtain streptavidin detectors. To optimize sensor design, we measured sensor response as a function of both surface linker length ($L=14,24,34$ or $44$ bp) and MB curtain length ($l_{MB}=$20, 40, or 60 bp).
@@
Figure 4: Lily pad sensors can quantify picomolar protein concentrations. SWV endpoint signals were measured for lily pads having 40 bp MB curtains and $L=$ 34 bp linkers using streptavidin concentrations from 1 to 1000 pM. Off- and on-signals were measured before and after each sample was incubated on a lily pad chip (one hour, ${}^{\circ}$). Mean and standard deviation were calculated from five measurements for each concentration. Signal changes in the interval $\mu_{0}\pm 3\sigma_{0}$ are shaded gray, where $\mu_{0}$ is the mean of the blank and $\sigma_{0}$ its standard deviation. LoD $<1$ pM.
Figure 5: Lily pad sensors with tightly bound analytes can be regenerated multiple times. (A) Schematic shows a lily pad being closed by streptavidin (i), reopened by ssDNA invaders via strand displacement (ii), and regenerated by the addition of new biotin-modified adaptors (iii). (B) Sensorgram demonstrating two rounds of streptavidin detection where notations are made for buffer washes (*) and the addition of 500 pM streptavidin (i), 100 nM invaders (ii), and 500 pM biotinylated adaptor strands. (C) Changes to SWV signal through four rounds of lily pad regeneration. In each round, streptavidin solution (10 $\mu$L, 500 pM) was reacted for one hour at ${}^{\circ}$ before a measurement was taken (high values). Lily pads with 40 bp MB reporters and $L=34$ bp were used for (B) and (C).
@@
To further demonstrate the modularity of the lily pad sensor and to test its capability for detecting larger analytes, DNA adapter strands were designed to display an aptamer that binds PDGF-BB (Fig. 6). Since the design of the sensor requires two binding events, one to the origami and one to the surface, PDGF-BB was chosen as the analyte due to its homodimeric nature, which allows for the same aptamer to be used on both the origami and the surface. Surface preparation and origami folding were performed as for previous designs, and adapter strands were added after origami had been tethered to the surface, as with the streptavidin sensor. The two adapters (Fig. 6A) were designed to have sequences ${}^{\prime}$ and ${}^{\prime}$, followed by a previously reported 36 nt aptamer with an apparent binding affinity to PDGF-BB of 36 fM(87). The resulting sensor achieved detection of PDGF-BB at concentrations as low as 500 pM (Fig. 6B), with a signal change of $20\%$. Sensor signals were stable after PDGF-BB addition for over 100 mins of serial monitoring every 5 min, highlighting the functional stability of lily pad sensors. The lowest concentration measured is within one order of magnitude of the best sensitivity achieved for previously reported aptamer-based PDGF-BB sensors(88). However, we note that because the goal of these experiments was simply to demonstrate platform modularity, no efforts were made to co-optimize the aptamer and sensor (e.g. matching aptamer size and MB curtain length); there is therefore, potential to improve the sensitivity of the PDGF-BB version of the lily pad sensor in future work.
@@
[…]
@@
For synthesis of double-stranded DNA linkers with defined length and single-stranded overhangs with specific sequences, we performed two sequential PCRs, extension and autosticky PCRs(98), using Lambda DNA purchased from Promega (WI) and Taq polymerase from New England Biolabs (MA). In the extension PCR, each of the DNA primers (sequences in Table S2), purchased from IDT (CA), consists of a 20 nt long template-binding region that determines the length of the linker and another 20 nt long extension region that determines the terminal sequences of the amplicon added to the sequence from the template DNA (Fig. S2, top). The amplicons from the first PCR, after purification by agarose gel extraction using Zymoclean Gel DNA Recovery Kits purchased from Zymo Research (CA), were used in the second autosticky PCR as template DNA. Each of the primers used in the autosticky PCR (Fig. S2, bottom) contains two domains separated by an abasic site, a 20 nt long template-binding region, identical to the extended sequence in the first PCR, and an overhang sequence chosen for binding to either the flat DNA origami or gold surface. The PCR products were purified using the DNA Clean and Concentrator Kit from Zymo Research and stored in 10 mM Tris buffer (pH 7.4) at ${}^{\circ}$.
@@
The DNA origami we used to prepare the lily pads is derived from a structure used in earlier work (49), which was designed in caDNAno(99) to be a flat, circular shape with a square opening. In that work the square opening gave the design a specific orientation on a microfabricated surface; here the square opening is irrelevant. For the purposes of this work, the design was modified so that extensions of staples on the ${}^{\prime}$ end would all appear on the same side (Fig. S7). Most staple strands (see Table S5 for sequences) were ordered from IDT unpurified at 100 $\mu$M in water and stored at ${}^{\circ}$. To introduce an analyte binding site near the center for DNA sensing experiments, one staple strand was replaced by a new DNA oligo (IDT, PAGE-purified) that has a 14 nt long ${}^{\prime}$ extension (shown in cyan in Fig. S7). To create 20 bp MB curtains: (1) 70 out of 234 total staples were extended on their ${}^{\prime}$ ends with a common 20 nt long single stranded linker and (2) an MB-modified DNA strand (IDT, dual-HPLC purified) with a sequence complementary to the 20 nt linker was hybridized to all 70 extended staples.
When preparing the lily pads with longer MB curtain lengths (40 or 60 bp) for streptavidin sensing experiments, slightly different schemes were used, as shown at right in Fig. 3B. For 40 bp curtains: (1) a 40 nt MB-modified strand (IDT, Dual-HPLC purified) was hybridized to the standard 20 nt extension, and (2) a single 20 nt “cap” ssDNA strand was hybridized to the ${}^{\prime}$ end of the MB-modified strand to create a fully duplexed curtain. For 60 bp curtains, origami were functionalized with 40 nt extensions that hybridized the 40 nt MB-modified strand via the same 20 nt overlap used for 40 bp curtains. This meant that to create a fully duplexed curtain, both the 20 nt cap described above, a second 20 nt cap were used (where the second cap was hybridized to the section of the staple extension that was proximal to the origami.)
[…]
@@
[…]
Two thiolated DNA oligos were purchased from IDT, one for analyte binding (${}^{\prime}$${}^{\prime}$) and the other for origami-linker tethering on gold (${}^{\prime}$-CGTAAACCCAGCGTCTTCACCACGATGAATACTCCCACCG- ${}^{\prime}$). In separate tubes, we mixed 1 $\mu$L of 100 $\mu$M thiolated DNA oligos and 1 $\mu$L of 10 mM tris(2-carboxyethyl) phosphine hydrochloride (TCEP, Sigma Aldrich) and incubated for 1 hour at room temperature to reduce the disulfide bonds. Then the solutions were mixed and diluted to 100 nM of each DNA in 1$\times$ PBS buffer (pH 7.4). 20 $\mu$L of the solution was introduced into a silicone gasket well on a freshly prepared TS-gold chip and a Teflon cell (CH instrument, TX) was assembled creating about 1.5 mL of reaction volume that allows three electrode connections. After 1 hour incubation at room temperature, the solution was exchanged with 50 $\mu$L of 10 mM 6-mercapto-1-hexanol (6-MCH, Sigma Aldrich) in 1$\times$ PBS buffer followed by overnight incubation at room temperature for formation of a passivating layer on the gold surface. After this step, the surface presumably had a random distribution of analyte binding and origami linker strands in a 1:1 ratio, with the spaces between these oligos filled by 6-MCH.
@@
A Metrohm PGSTAT 128N (Netherlands) potentiostat was used for SWV (Fig. S5). After equilibrating the cell at -0.15 V, SWV measurements were performed at a frequency of 10 Hz (see Fig. S6 for optimization) with an amplitude of 25 mV between -0.15 V and -0.4 V relative to Ag/AgCl reference electrode in 1 mL of TAE/Mg at ${}^{\circ}$; temperature was controlled by using a Coy Labs glove box—without strict temperature control sensor response was too variable. SWV voltammograms were recorded every 5 mins for 1 hour before and 5 hours after the addition of analyte DNA, if not otherwise specified. The first hour of measurements before adding analyte is used to set the baseline, off-signal, and measure the fold-increase. To determine the peak current value a linear baseline is subtracted from the measurement and the MB peak is isolated. This peak is then fit to a Gaussian(100) and its maximum value is recorded as the peak current (Fig. S5).
@@
Two biotin-modified strands were ordered from IDT, one to create a surface binding site, complementary to the DNA analyte binding thiolated strand (5’-Biotin-CAGATATCAAAGCT-3’), and the other to create a binding site on origami, complementary to the DNA analyte binding tail (${}^{\prime}$${}^{\prime}$). The DNA sensing chip was incubated for 30 mins with 500 pM of each of these two strands in TAE/Mg at room temperature. This resulted in an $L=14$ sensor. After rinsing with TAE/Mg, SWV measurements were performed for 500 pM streptavidin (Thermo Fischer Scientific) at ${}^{\circ}$ in the same buffer.
To create streptavidin sensors with dsDNA surface linkers having a length greater than 14 bp (Fig. 3C), we immobilized one of three different ${}^{\prime}$-thiolated ssDNAs (10, 20, or 30 nt long) instead of the standard ${}^{\prime}$-thiolated sequence; we then performed a 6-MCH passivation step. Next, a 30 minute incubation was performed to simultaneously hybridize the lily pads to the surface, and add bridging strands where necessary. In particular, 5 $\mu$M of 24, 34, or 44 nt ssDNA bridging strands were added. The ${}^{\prime}$ sequence of these strands was complementary to one of the 10, 20, or 30 nt ssDNA already on the surface (as appropriate), so they formed dsDNA complexes proximal to the surface; the 14 nt ${}^{\prime}$ end of these strands projected ssDNA tails having the sequence ${}^{\prime}$ into solution. In the next step where the biotin adapter strands (500 pM) were added, one of the adapters bound to the ${}^{\prime}$ tails extending from the surface, forming dsDNA linkers with $L=$ 24, 34, or 44 bp.
For LoD experiments in Fig. 4, after rinsing off excess biotin-adaptor strands, we measured SWV signal five times before and five times after incubation with the relevant streptavidin solution on the chip; the pre- and post- binding signal for each chip was calculated as the mean of these values. Between the measurements, 10 $\mu$L of varying concentrations of streptavidin samples were introduced into a gasket well on the chip, and sealed with a piece of Parafilm to prevent sample evaporation. After 1 hour incubation at ${}^{\circ}$, chips were thoroughly washed with TAE/Mg buffer before the endpoint measurements. This process was repeated for 5 different chips per streptavidin concentration.
@@
To create PDGF-BB detecting lily pad sensors, DNA aptamers for PDGF-BB protein (${}^{\prime}$${}^{\prime}$) were ordered from IDT with two 14 nt extensions, one at ${}^{\prime}$-end for binding to the thiolated DNA strands immobilized on gold surface (${}^{\prime}$${}^{\prime}$) and the other at ${}^{\prime}$-end for the single stranded DNA tail on origami (${}^{\prime}$${}^{\prime}$) that were used for the DNA analyte and streptavidin detection studies (see Table S3). PDGF-BB protein was ordered from PeproTech, Inc. (NJ). After preparing a standard DNA analyte detecting lily pad sensor, the chip was incubated with 1 nM each of the PDGF-BB aptamers with extensions in TAE/Mg at room temperature. The chip, assembled in a Teflon cell, was rinsed in the same buffer and connected to potentiostat. SWV measurements were performed as described above in 1 mL of TAE/Mg at ${}^{\circ}$. After recording SWV voltammograms every 5 mins for 1 hour, 0, 500 pM, 1 nM, or 2 nM of PDGF-BB protein in the same buffer was added.

### v2

@@ Abstract
This manuscript was compiled on June 24, 2024
@@
Figure 1: Lily pad sensors can be used for the electrochemical detection of biological analytes, here a DNA single strand. (A) A flat, disk-shaped DNA origami with a square hole (49) carrying a long dsDNA linker is assembled from a mixture of long ssDNA scaffolds, ssDNA staples, 70 staples with 5-extensions for MB reporters and for analyte binding, 20 nt MB-modified DNA oligos, and long dsDNA linkers with ssDNA overhangs at both ends. Inset shows AFM of a lily pad. (B) Two types of thiol-modified ssDNA are immobilized on a template-stripped gold surface: one for tethering the lily pad origami via hybridization with one of the linker overhang sequences, and the other for analyte binding to the gold surface. Closing events occur when a DNA analyte binds both to the binding site on the origami and the binding site on the surface, enabling electron transfer between MB reporters and the gold electrode. Inset shows an exemplary square wave voltammogram of a sensor in the open state (left), and a voltammogram in the closed state after DNA analyte addition (right). The peak current is calculated as the heights of the peaks relative to the underlying baseline.
@@ Results
Figure 1 shows the design of our lily pad device. The device consists of a 100 nm diameter disk-shaped DNA origami (see methods) that is attached to a long dsDNA linker (Fig. 1A) which, depending on experiment, ranges from 300 to 3000 base pairs (bp) in length. The distal end of the linker is attached to a gold electrode through DNA hybridization: the distal end of the linker displays a single-stranded DNA (ssDNA) overhang, and binds to a complementary ssDNA that is immobilized on the gold electrode. E-DNA/E-AB sensors often use roughened gold electrodes to maximize the surface area for their small ssDNA sensing molecules(56, 57), despite the resultant structural heterogeneity of such electrodes. Here, in contrast, we use ultraflat, template-stripped gold chips(58) to minimize the distance between the entire bottom face of the origami disk and the electrode surface upon analyte binding. Our lily pad sensors use 70 MB molecules as reporters to transfer electrons to the electrode; they are attached to one side of each origami via 20 bp-long ssDNAs hybridized to overhangs on 70 modified staple strands (Fig. 1B)—this forms a curtain of MB-modified DNA strands hanging from the origami. Detection of a ssDNA analyte sequence xy is achieved through binding to two DNA sequences, x and y, which are complementary to subsequences x and y of xy, respectively. Sequence x is positioned on the origami as an extension to a central staple, and a thiol-modified version of y is immobilized on the gold electrode surface via an Au-S bond. When both x and y in a DNA analyte bind to x and y on the lily pad and the underlying electrode, the conformation of the lily pad changes from the “open” to the “closed” state. Each closing event brings the 70 MB curtain into proximity with the gold surface, facilitating electron transfer between MB and electrode and increasing voltammetric peak currents (Fig. 1B, inset).
@@
Figure 2: Lily pad closing events can be monitored via voltammetry in real-time. (A) 500 pM of each of two different ssDNA analytes, with 28-bp sequences xy (blue) and yx (orange), were added at time $t=0$ at 34C. The difference in kinetics for xy and yx binding can be understood in terms of difference spacing between the origami and gold surface in their respective bound states (bottom left). For yx to bind (orange box), the origami must get much closer to the surface than in the case of xy (blue box), incurring an increased entropy penalty. Binding of yx may also be sterically hindered by the 20 bp long MB reporters that comprise a curtain projecting down from the origami (top left). (B) The length of the linker tethering the lily pad to the surface has little effect on sensor response: linker lengths of 332, 1037, and 2983 bp all show comparable behavior. (C) The functional form of sensor response varies with the concentration of ssDNA analyte. At analyte concentrations up to 500 picomolar (yellow diamonds), sensor response increases monotonically. At nanomolar concentrations (green circles), the system shows saturation effects, and the current decreases after one hour, potentially due to excess free analyte in solution reopening previously-closed lily pads via DNA strand displacement. ssDNA analyte xy added at $t=0$.
@@
Because of the lily pad’s modularity, it is trivial to modify into a detector for multimeric proteins with multiple identical binding sites. As a proof-of-principle (Fig. 3), we prepared lily pad sensors for detection of a model multimeric protein, streptavidin. We achieved this by simply adding two biotinylated adaptor strands to our basic DNA-sensing chips (Fig. 3A): the first adaptor, sequence x, was complementary to the 14 nt ssDNA tail (x) on the origami; the second, sequence y, was complementary to the 14 nt oligos (y) on the surface. Thus chips with the lily pads used for ssDNA detection were incubated with the biotin-modified adaptors to obtain streptavidin detectors. To optimize sensor design, we measured sensor response as a function of both surface linker length ($L=14,24,34$ or $44$ bp) and MB curtain length ($l_{MB}=$20, 40, or 60 bp).
@@
Figure 4: Lily pad sensors can quantify picomolar protein concentrations. SWV endpoint signals were measured for lily pads having 40 bp MB curtains and $L=$ 34 bp linkers using streptavidin concentrations from 1 to 1000 pM. Off- and on-signals were measured before and after each sample was incubated on a lily pad chip (one hour, 34C). Mean and standard deviation were calculated from five measurements for each concentration. Signal changes in the interval $\mu_{0}\pm 3\sigma_{0}$ are shaded gray, where $\mu_{0}$ is the mean of the blank and $\sigma_{0}$ its standard deviation. LoD $<1$ pM.
Figure 5: Lily pad sensors with tightly bound analytes can be regenerated multiple times. (A) Schematic shows a lily pad being closed by streptavidin (i), reopened by ssDNA invaders via strand displacement (ii), and regenerated by the addition of new biotin-modified adaptors (iii). (B) Sensorgram demonstrating two rounds of streptavidin detection where notations are made for buffer washes (*) and the addition of 500 pM streptavidin (i), 100 nM invaders (ii), and 500 pM biotinylated adaptor strands. (C) Changes to SWV signal through four rounds of lily pad regeneration. In each round, streptavidin solution (10 $\mu$L, 500 pM) was reacted for one hour at 34C before a measurement was taken (high values). Lily pads with 40 bp MB reporters and $L=34$ bp were used for (B) and (C).
@@
To further demonstrate the modularity of the lily pad sensor and to test its capability for detecting larger analytes, DNA adapter strands were designed to display an aptamer that binds PDGF-BB (Fig. 6). Since the design of the sensor requires two binding events, one to the origami and one to the surface, PDGF-BB was chosen as the analyte due to its homodimeric nature, which allows for the same aptamer to be used on both the origami and the surface. Surface preparation and origami folding were performed as for previous designs, and adapter strands were added after origami had been tethered to the surface, as with the streptavidin sensor. The two adapters (Fig. 6A) were designed to have sequences x and y, followed by a previously reported 36 nt aptamer with an apparent binding affinity to PDGF-BB of 36 fM(87). The resulting sensor achieved detection of PDGF-BB at concentrations as low as 500 pM (Fig. 6B), with a signal change of $20\%$. Sensor signals were stable after PDGF-BB addition for over 100 mins of serial monitoring every 5 min, highlighting the functional stability of lily pad sensors. The lowest concentration measured is within one order of magnitude of the best sensitivity achieved for previously reported aptamer-based PDGF-BB sensors(88). However, we note that because the goal of these experiments was simply to demonstrate platform modularity, no efforts were made to co-optimize the aptamer and sensor (e.g. matching aptamer size and MB curtain length); there is therefore, potential to improve the sensitivity of the PDGF-BB version of the lily pad sensor in future work.
@@
Similar to E-AB sensors, our platform is reagentless and signaling relies on target-binding induced conformational change—it thus has the potential to work in complex biological matrices, in vivo and ultimately in awake behaving animals. Achieving this potential will require strategies to increase the lifespan of the electrode and monolayer in serum(96). Hardening origami to nuclease degradation, low Mg2+ concentrations, and nonspecific binding in vivo, while maintaining its structure and functionality, would also be necessary. Such hardening has been achieved by a variety of methods(97) including increasing helical packing density, chemical cross-linking, coating with block copolymers and functionalization with unnatural nucleotides and end-groups; these methods should translate readily to the lily pad.
@@
For synthesis of double-stranded DNA linkers with defined length and single-stranded overhangs with specific sequences, we performed two sequential PCRs, extension and autosticky PCRs(98), using Lambda DNA purchased from Promega (WI) and Taq polymerase from New England Biolabs (MA). In the extension PCR, each of the DNA primers (sequences in Table S2), purchased from IDT (CA), consists of a 20 nt long template-binding region that determines the length of the linker and another 20 nt long extension region that determines the terminal sequences of the amplicon added to the sequence from the template DNA (Fig. S2, top). The amplicons from the first PCR, after purification by agarose gel extraction using Zymoclean Gel DNA Recovery Kits purchased from Zymo Research (CA), were used in the second autosticky PCR as template DNA. Each of the primers used in the autosticky PCR (Fig. S2, bottom) contains two domains separated by an abasic site, a 20 nt long template-binding region, identical to the extended sequence in the first PCR, and an overhang sequence chosen for binding to either the flat DNA origami or gold surface. The PCR products were purified using the DNA Clean and Concentrator Kit from Zymo Research and stored in 10 mM Tris buffer (pH 7.4) at -20C.
@@
The DNA origami we used to prepare the lily pads is derived from a structure used in earlier work (49), which was designed in caDNAno(99) to be a flat, circular shape with a square opening. In that work the square opening gave the design a specific orientation on a microfabricated surface; here the square opening is irrelevant. For the purposes of this work, the design was modified so that extensions of staples on the 5 end would all appear on the same side (Fig. S7). Most staple strands (see Table S5 for sequences) were ordered from IDT unpurified at 100 $\mu$M in water and stored at -20C. To introduce an analyte binding site near the center for DNA sensing experiments, one staple strand was replaced by a new DNA oligo (IDT, PAGE-purified) that has a 14 nt long 5 extension (shown in cyan in Fig. S7). To create 20 bp MB curtains: (1) 70 out of 234 total staples were extended on their 5 ends with a common 20 nt long single stranded linker and (2) an MB-modified DNA strand (IDT, dual-HPLC purified) with a sequence complementary to the 20 nt linker was hybridized to all 70 extended staples.
When preparing the lily pads with longer MB curtain lengths (40 or 60 bp) for streptavidin sensing experiments, slightly different schemes were used, as shown at right in Fig. 3B. For 40 bp curtains: (1) a 40 nt MB-modified strand (IDT, Dual-HPLC purified) was hybridized to the standard 20 nt extension, and (2) a single 20 nt “cap” ssDNA strand was hybridized to the 3 end of the MB-modified strand to create a fully duplexed curtain. For 60 bp curtains, origami were functionalized with 40 nt extensions that hybridized the 40 nt MB-modified strand via the same 20 nt overlap used for 40 bp curtains. This meant that to create a fully duplexed curtain, both the 20 nt cap described above, a second 20 nt cap were used (where the second cap was hybridized to the section of the staple extension that was proximal to the origami.)
To synthesize origami, we mixed 8094 nt long scaffold strands (p8094 from Tilibit, Germany), with the staple strand mixture, MB-modified DNA oligos (IDT), and the dsDNA linker from PCRs to the final concentrations given in Table S4 (1:1 scaffold:linker, 5 nM each). One staple on the left side of the origami (yellow orange in Fig. S7) was omitted to leave a position on the scaffold at which the dsDNA linker overhang could bind. Throughout this work, we refer to 1$\times$ TAE buffer (Biorad, pH 7.5) with 12.5 mM MgCl2 as “TAE/Mg”. 100 $\mu$L of scaffold/staple/linker mixture in TAE/Mg was heated to 90C for 5 min and annealed from 90C to 20C at -1C/min. The final concentration of origami, based on the initial scaffold concentration was 5 nM and the solution was then diluted for use to 2 nM in TAE/Mg buffer.
@@
We use the template-stripping (TS) method to prepare ultraflat (see AFM, Fig. S4) gold surfaces as a substrate for lily pad sensors(58). First, a 200 nm gold film was deposited on a 4 inch silicon wafer (University Wafers, MA) using a Labline electron beam evaporator (Kurt J. Lesker Company, PA) at the Kavli Nanoscience Institute Lab at Caltech. The wafer was then cut to $5\times 8$ mm2 chips by Dynatex GST-150 Scriber/Breaker. A $10\times 10$ mm2 glass coverslip (#2) was rinsed with acetone, isopropyl alcohol, and water, blown dry with nitrogen, and cleaned by oxygen plasma for 5 min in a PE-50 plasma system (Plasma Etch, NV). About 1 $\mu$L of UV-curable adhesive (Noland, No.61) was applied on a clean glass and a gold chip was placed on top of the adhesive. The adhesive was then cured via a long-wave UV irradiation for 1 hour (66). Right before use, the silicon wafer was pried off the gold/adhesive/glass layers using a razor blade, exposing the ultraflat gold surface on a glass coverslip. To create a well and isolate the reactive area, a silicone gasket (Grace Bio-Labs Press-To-Seal silicone isolator, 2 mm diameter) was glued on top and copper tape was used to form an electrical connection.
Two thiolated DNA oligos were purchased from IDT, one for analyte binding (5-HS-TTTTTAGCTTTGATATCTG-3) and the other for origami-linker tethering on gold (5-CGTAAACCCAGCGTCTTCACCACGATGAATACTCCCACCG- TTTTT-SH-3). In separate tubes, we mixed 1 $\mu$L of 100 $\mu$M thiolated DNA oligos and 1 $\mu$L of 10 mM tris(2-carboxyethyl) phosphine hydrochloride (TCEP, Sigma Aldrich) and incubated for 1 hour at room temperature to reduce the disulfide bonds. Then the solutions were mixed and diluted to 100 nM of each DNA in 1$\times$ PBS buffer (pH 7.4). 20 $\mu$L of the solution was introduced into a silicone gasket well on a freshly prepared TS-gold chip and a Teflon cell (CH instrument, TX) was assembled creating about 1.5 mL of reaction volume that allows three electrode connections. After 1 hour incubation at room temperature, the solution was exchanged with 50 $\mu$L of 10 mM 6-mercapto-1-hexanol (6-MCH, Sigma Aldrich) in 1$\times$ PBS buffer followed by overnight incubation at room temperature for formation of a passivating layer on the gold surface. After this step, the surface presumably had a random distribution of analyte binding and origami linker strands in a 1:1 ratio, with the spaces between these oligos filled by 6-MCH.
@@
A Metrohm PGSTAT 128N (Netherlands) potentiostat was used for SWV (Fig. S5). After equilibrating the cell at -0.15 V, SWV measurements were performed at a frequency of 10 Hz (see Fig. S6 for optimization) with an amplitude of 25 mV between -0.15 V and -0.4 V relative to Ag/AgCl reference electrode in 1 mL of TAE/Mg at 34C; temperature was controlled by using a Coy Labs glove box—without strict temperature control sensor response was too variable. SWV voltammograms were recorded every 5 mins for 1 hour before and 5 hours after the addition of analyte DNA, if not otherwise specified. The first hour of measurements before adding analyte is used to set the baseline, off-signal, and measure the fold-increase. To determine the peak current value a linear baseline is subtracted from the measurement and the MB peak is isolated. This peak is then fit to a Gaussian(100) and its maximum value is recorded as the peak current (Fig. S5).
@@
Two biotin-modified strands were ordered from IDT, one to create a surface binding site, complementary to the DNA analyte binding thiolated strand (5’-Biotin-CAGATATCAAAGCT-3’), and the other to create a binding site on origami, complementary to the DNA analyte binding tail (5-CTGAATGGTACGGA-Biotin-3). The DNA sensing chip was incubated for 30 mins with 500 pM of each of these two strands in TAE/Mg at room temperature. This resulted in an $L=14$ sensor. After rinsing with TAE/Mg, SWV measurements were performed for 500 pM streptavidin (Thermo Fischer Scientific) at 34C in the same buffer.
To create streptavidin sensors with dsDNA surface linkers having a length greater than 14 bp (Fig. 3C), we immobilized one of three different 3-thiolated ssDNAs (10, 20, or 30 nt long) instead of the standard 5-thiolated sequence; we then performed a 6-MCH passivation step. Next, a 30 minute incubation was performed to simultaneously hybridize the lily pads to the surface, and add bridging strands where necessary. In particular, 5 $\mu$M of 24, 34, or 44 nt ssDNA bridging strands were added. The 5 sequence of these strands was complementary to one of the 10, 20, or 30 nt ssDNA already on the surface (as appropriate), so they formed dsDNA complexes proximal to the surface; the 14 nt 3 end of these strands projected ssDNA tails having the sequence y into solution. In the next step where the biotin adapter strands (500 pM) were added, one of the adapters bound to the y tails extending from the surface, forming dsDNA linkers with $L=$ 24, 34, or 44 bp.
For LoD experiments in Fig. 4, after rinsing off excess biotin-adaptor strands, we measured SWV signal five times before and five times after incubation with the relevant streptavidin solution on the chip; the pre- and post- binding signal for each chip was calculated as the mean of these values. Between the measurements, 10 $\mu$L of varying concentrations of streptavidin samples were introduced into a gasket well on the chip, and sealed with a piece of Parafilm to prevent sample evaporation. After 1 hour incubation at 34C, chips were thoroughly washed with TAE/Mg buffer before the endpoint measurements. This process was repeated for 5 different chips per streptavidin concentration.
@@
To create PDGF-BB detecting lily pad sensors, DNA aptamers for PDGF-BB protein (5-CAGGCTACGGCACGTAGAGCATCACCATGATCCTG-3) were ordered from IDT with two 14 nt extensions, one at 3-end for binding to the thiolated DNA strands immobilized on gold surface (5-CAGATATCAAAGCT-3) and the other at 5-end for the single stranded DNA tail on origami (5-CTGAATGGTACGGA-3) that were used for the DNA analyte and streptavidin detection studies (see Table S3). PDGF-BB protein was ordered from PeproTech, Inc. (NJ). After preparing a standard DNA analyte detecting lily pad sensor, the chip was incubated with 1 nM each of the PDGF-BB aptamers with extensions in TAE/Mg at room temperature. The chip, assembled in a Teflon cell, was rinsed in the same buffer and connected to potentiostat. SWV measurements were performed as described above in 1 mL of TAE/Mg at 30C. After recording SWV voltammograms every 5 mins for 1 hour, 0, 500 pM, 1 nM, or 2 nM of PDGF-BB protein in the same buffer was added.